OpenStudy (anonymous):

Find the equation of the line passing through the points (-1, 9), (3, -7).

OpenStudy (mitu12):

use y1-y2\x2-x1

OpenStudy (anonymous):

\[(x1-y1)(x2-y2)\] \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

umm ??

OpenStudy (anonymous):

I dont know how to figure which one is y1 or x2

OpenStudy (mitu12):

9-(-7)=?

OpenStudy (mitu12):

the first number is x and the second number is y

OpenStudy (anonymous):

16

OpenStudy (mitu12):

yes

OpenStudy (mitu12):

-1-3=?

OpenStudy (anonymous):

my options are y = 4x + 5 y = -4x + 5 y = -4x + 13 y = -4x - 5

OpenStudy (anonymous):

-4

OpenStudy (mitu12):

yes

OpenStudy (anonymous):

so which one would it be?

OpenStudy (mitu12):

are you sure those are the options

OpenStudy (anonymous):

yes

OpenStudy (mitu12):

one moment

OpenStudy (anonymous):

ok

OpenStudy (mitu12):

it will be B

OpenStudy (anonymous):

ok thanks

OpenStudy (mitu12):

no problem